OpenStudy (anonymous):

Which of these statements are true? 1 km = 10^3 m 10^3 ms = 10^9 ns 10^3 m = 10^6 mm 1 km = 10^5 mm 1 cm^3 = 10 mL

OpenStudy (anonymous):

10^3 m = 10^6 mm and 1 km = 10^3 m are true

OpenStudy (anonymous):

what about ms and ns

OpenStudy (anonymous):

millisecond and nanoseconds